OpenStudy (anonymous):

heeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeeellllllllllllllllllllllllllllllllllllllllllllllllllllpppppppppppppppppppppppppppppppppppppppppppppp . Madeline works at a home improvemt store .a formula for a certain color calls for 3/5 of a liter of red paint . Madeline wants to make 2/3 less paint than the formula creates . How much red paint should Madeline use ?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@Callisto

OpenStudy (anonymous):

@Compassionate

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@mayankdevnani

OpenStudy (anonymous):

@Jhannybean

OpenStudy (michele_laino):

I think the answer is: \[\frac{ 3 }{ 5 }-(\frac{ 3 }{ 5 }\times \frac{ 2 }{ 3 })\]